if the average speed of an orbiting space shuttle is 19,800mi/h, determine the time required for it to circle earth
1
Expert's answer
2019-09-05T09:09:19-0400

"r=3963\ miles"

"h=200\ miles"

"R=r+h"


"T=\frac{2\pi R}{v}"

"T=\frac{2\pi (3963+200)}{19800}=1.32\ h"